OpenStudy (tennistar):

Find the exact value of the radical expression in simplest form. 2 the square root of 7 times x + the square root of 7 times x - 3 the square root of 7 times x - 7

OpenStudy (tennistar):

Answer choices -7 4 the square root of 7 times x - 7 the square root of 7 times x 3 the square root of 7 times x - 7

OpenStudy (tennistar):

\[2\sqrt{7x}+\sqrt{7x}-3\sqrt{7x}-7\]

OpenStudy (tennistar):

@ash2326

OpenStudy (tennistar):

I dont understand this

OpenStudy (ash2326):

@Tennistar it's quiet easy if you have \[2x+x-3x\] How would you simplify this?

OpenStudy (ash2326):

* quite sorry

OpenStudy (tennistar):

ummm I dont know

OpenStudy (ash2326):

ok, what's \[2+1-3\] ???

OpenStudy (tennistar):

0

OpenStudy (ash2326):

Good, so if we have \[2x+x-3x\] We can notice all the terms have x here, so we'll simplify it using normal arithmetic just like you did for (2+1-3) So what would get here? @Tennistar

OpenStudy (tennistar):

0 right

OpenStudy (ash2326):

Yeah , if we have \[2x+x-x\] What would you get?

OpenStudy (tennistar):

because you remove the x from all parts by dividing

OpenStudy (tennistar):

can you help me solve it step by step

OpenStudy (ash2326):

Yeah we divide all the terms by x and place it outside the bracket When you solved \[2x+x-3x\] You actually did \[x(2+1-3)\] \[x\times 0=>0\]

OpenStudy (ash2326):

Yeah, I'll help you Did you understand this?

OpenStudy (tennistar):

I really need to know and yes

OpenStudy (ash2326):

okay here we have \[2\sqrt{7x}+\sqrt{7x}-3\sqrt{7x}-7\] Notice the first three terms \[\underline{2\sqrt{7x}+\sqrt{7x}-3\sqrt{7x}}-7\] First simplify them, just like you did for \((2x+x-3x\)) try this

OpenStudy (tennistar):

so divide by 7x*sqrt all parts?

OpenStudy (ash2326):

Yeah, divide and take it outside and then simplify the numbers

OpenStudy (tennistar):

adn have 2 + 1-3-7

OpenStudy (tennistar):

-7

OpenStudy (tennistar):

where?

OpenStudy (ash2326):

Sorry, yeah, you're right. I'll show you the step wise approach

OpenStudy (ash2326):

\[2\sqrt{7x}+\sqrt{7x}-3\sqrt{7x}+7\] Take \(\sqrt{7x}\) common from the first three terms \[\sqrt{7x}(2+1-3)-7\] \[\sqrt{7x}\times 0-7=>-7\] voila

OpenStudy (ash2326):

Do you get this?

OpenStudy (tennistar):

yes

OpenStudy (ash2326):

@Tennistar I really liked that you are so eager to learn. good work and way to go :)